# Photoluminescence from metal nanostructures - dependence on size

### v1

@@ 3.1.2 Determination of the PL
As already noted above, the expression for PL (9) can be further simplified for a sphere. As discussed in Section 3.1.1, the temperature is fairly uniform in the sphere and the space-dependence of in Eq. (9) originates only from that of , as indicated by Eq. (2.1). Further, we rely on the calculations of the mean free path of the non-thermal electrons 30 to justify the neglect of spatial broadening of the non-thermal electron distribution. These results motivate the use of the analytic result (2.1) also for non-uniform fields, i.e., to approximate the level of non-thermal electron population by simply summing over its local value across the particle volume. Accordingly, (5) can be expressed in terms of the spatially averaged, . With this consideration, (2.1) becomes independent of position. The volume integration over is now replaceable by the absorption cross-section of the sphere, which in turn, can be expressed as . Thus, Eq. (9) simplifies to
@@
Using this expression, we can determine the total PL solely via the absorption cross-section density evaluated at the pump frequency (, Fig. 2(b)) and emission frequency (, Fig. 10(a)), as well as the electron temperature (Fig. 2(c)). As seen in Fig. 3(a)-(c), the qualitative behaviour of the PL (Eq. (9)) and its approximation (Eq. (12)) is nearly identical except for the former being slightly higher for nm.
@@
While the computational simplification associated with Eq. (12) is minor, its more significant contribution is to allow us to pinpoint the origins of the behaviour observed in Fig. 3. Specifically, the PL in Fig. 3(a)-(c) pertains roughly to room temperature for all sphere sizes making the Bose function () essentially size-independent. In this scenario, the PL varies with sphere size as the product of (at the absorption wavelength), (at the emission wavelength) and the volume, . For spheres of radius nm (i.e., safely within the quasi-static regime), has a very weak dependence on the volume (see Fig. 9) so that the PL scales as . For larger spheres, the sub-volume scaling of the absorption and temperature, along with variations in with particle size owing to the sphere’s modal response (see Ref. [ 35] and Fig. 2(b)) result in the observed deviation from the volume scaling, namely, the peaks and troughs in the PL shown in all subplots of Fig. 3. Specifically, the non-monotonic behaviour in the PL observed for the larger sizes are due to the combined effect of the system tuning into and then out of the dipole resonance, and then tuning into the quadrupole mode. In that respect, considering the PL as being a volume effect for small spheres, and as a surface effect for larger spheres, is a crude description, which misses the dominant effect of the modal structure.
@@
Fig. 3(d)-(f) shows the PL for a higher intensity which results in a size-dependent temperature, varying from room temperature to (see Fig. 2(c)). The higher and varying temperature affects differently the different parts of the PL spectrum. First, it hardly affects the SE (Fig. 3(e)-(f)) which thus follows the same trend observed for the low-intensities, at least up to temperatures for which thermo-optic effects are still negligible 38, 62 (see Section 3.1.3). Indeed, for SE, the non-thermal contribution from the PA term dominates (see Eq. (2.1), Fig. 1 and Fig. 11), which is nearly temperature-independent since it originates from the non-thermal electron shoulders given in Eq. (5) (see SI Section S3 of Ref. [ 11]). In contrast, the aSE is exponentially sensitive to the temperature via the Bose function (see Fig. 1). As a result, the size-dependence of the PL is much stronger (see Fig. 3(d)), exhibiting superlinear scaling with the volume.
@@
Deeper into the aSE regime, the PA term becomes stronger than the 1PA term, see Fig. 1. To demonstrate this, we calculate the PL for a longer (off-resonant) excitation wavelength ( nm, i.e., smaller ), see Fig. 12. In this case, while the near aSE (Fig. 12(b) and (f)) grows as , for frequencies sufficiently deep into the aSE, the PL scales as at low intensity excitation (Fig. 12(a)) and as at high intensity illumination (Fig. 12(d)).
@@
Finally, in order to show the generality of our approach, and to demonstrate that the behaviour we identify is not qualitatively affected by the choice of the exciting laser frequency (e.g., its position with respect to resonance), we replicate the results of Fig. 3 (for which the illumination is on the red-side of the plasmon resonance) in Fig. 13 and Fig. 14 for backgrounds with higher refractive indices. This results in a red-shift of the resonances and introduces the higher-order modal response to the PL emission, particularly, for the larger spheres. However, the qualitative behaviour is similar to that with the lower background refractive index.
@@ 3.1.3 The thermo-optic effect
Illumination at high intensity can induce significant changes in the temperature of the spheres, consequently altering their permittivity via the so-called thermo-optic effect 63, 37, 64, and as a result, modify the absorptivity (hence, the emissivity) and finally the PL. To analyze this effect, we assume that the permittivity of the metal and the thermal conductivity of the surrounding medium depend linearly on the change of temperature, compute the thermo-derivative of the metal permittivity from ellipsometry data 65 and set the thermo-derivative of the thermal conductivity of the surrounding medium to Wm-1K-2, as in Ref. [ 62]; the temperature of the nanoparticle is then calculated using the method described by Un et al. 62.
Overall, as the intensity (hence, temperature) increases, the real part of the permittivity becomes more negative, while the imaginary part increases to higher positive values at most frequencies. This results in a blue shift of the resonance peak and a broadening of its spectral width, leading to enhanced absorption at off-resonance frequencies, see Fig. 15 and Ref. [ 66, 67, 65, 37, 64]. As a result, slightly decreases at nm and slightly increases at nm, as seen in Fig. 2(a). The decrease in the former case leads to a reduction in the temperature for the hottest spheres (around nm), and hence, to the reduction of the aSE observed in Fig. 3(d). However, the SE PL is hardly affected by the temperature. Thus, the increase in at longer wavelengths results in a slight increase of the PL, as observed in Fig. 3(e) and (f). For excitation at nm, the temperature of the spheres (especially for nm) increases slightly compared to the value calculated for the linear case, leading to a slight increase in the PL for aSE, as shown in Fig. 12(d)-(e). The thermo-optic increase of and leads to a slight increase in SE, as observed in Fig. 12(f).
3.1.4 The quantum size effect
It is well known that the imaginary part of metal permittivites are found to be higher when the dimension of the metal structures falls below 10-20 nm 68. This effect is usually modelled by inserting a correction to the collision rate appearing in the Drude permittivity which grows inversely with the structure size 69. In that respect, the effect of the small size on the PL is similar to that of the thermo-optic effect, but is limited mostly to the small particle sizes, for which the PL is anyhow weaker. For this reason, we do not show a specific numeric example of this effect for spheres.
@@ 3.1.5 Comparison to single particle experiments
Our theoretical predictions can be compared to various single particle PL measurements. First, Fig. 4(a) shows that our calculations align with the experimental observations of the normalized SE PL (spectrally integrated over the range of nm to nm) measured from single gold nanospheres of different sizes, as reported by Gaiduk et al. (see Fig. S5 of Ref. [ 24]). The illumination intensity was adjusted to keep the temperature rise relatively low (approximately below K). The PL is calculated using Eq. (9) with the refractive index of gold obtained from Ref. [ 70], and those of glass and glycerol being and , respectively. The normalized PL roughly scales with the volume until nm, but seems to saturate upon further increase in size, in line with our prediction.
@@
Size-dependence of the PL from larger particles was studied by Hu et al. 25 who demonstrated that the spectral peak of the PL from gold nano-disks on a SiO2-coated absorptive (silicon) substrate initially increases, reaches a maximum, and then decreases significantly with further increases in size The absorption cross-section densities and the of the disks are calculated using COMSOL, considering normal incidence and emission perpendicular to the substrate, with the permittivity data for gold and Si taken from Ref. [ 70] and Ref. [ 71], respectively. Calculations based on Eq. (9) exhibit a qualitatively similar trend (even if at a slight spectral shift).
@@
Lastly, our theoretical prediction of PL are compared with the PL measurements by Lin et al. 27. This work showed the spectrally-integrated PL from single nanorods of different lengths with constant diameter along their main axis. In contrast to the nanospheres (and in similarity to layers, shown later), the PL monotonically decreases with nanorod length. These results are reproduced in Fig. 4(c) along with the theoretical prediction, i.e., the spectral integration of Eq. (9). The optical response is calculated in COMSOL, where the substrate refractive index is and the permittivity of the silver rods is taken from Ref. [ 70]. The temperature is estimated with an approximate expression 60 and remains close to the ambient temperature (0.1% variation) for the laser illumination intensity considered ( mW/cm2). Good qualitative agreement is obtained for nanorods longer than nm, however, our prediction underestimates the PL for shorter nanorods, thus providing only a qualitative match for this regime (specifically, in a 30 nm wide wavelength range close to the pump). A similar observation was recently reported above the pump wavelength 36, necessitating a phenomenological adjustment of the theoretical prediction.
@@
Remarkably, while our theory (9) includes only emission events occurring in the conduction band, it successfully predicts the dependence of the emission on the particle size even in spectral regimes where interband emission events should occur (specifically, for the spheres and nano-disks (Fig. 4(a)-(b), respectively). This implies that, in these cases, interband transitions modify the emission features in a modest quantitative manner, but not qualitatively. We expect that inclusion of such transitions may yield a better quantitative match to the experimental results.
3.2 Layers under continuous illumination
@@
We now turn to study the PL from a thin layer of a Drude metal using the approach used for the spheres. Specifically, we examine a thin silver layer on a glass substrate illuminated at normal incidence (hence, no dependence on polarization , and , where is the component of k normal to the surface of the layer) by a focused Gaussian CW beam with an intensity profile as depicted in Fig. 5(a).
@@ 3.2.1 Determination of the electron temperature
The absorbed power density and temperature distribution calculated by solving the heat equation (10) using the electromagnetic heat module of COMSOL Multiphysics, are presented in Figs. 5(b)-(c), respectively. The absorbed power density decreases exponentially along the illumination direction (-axis) due to the short penetration depth (Fig. 5(d)). However, along the radius it exhibits nearly the same distribution as that of the incident intensity, as shown in Fig. 5(e).
@@
Generically, is uniform in the plane of the film (i.e., it is independent of and ) and varies only along the thickness of the film (i.e., it is only -dependent). The average value of along the thickness can be derived from the experimentally measurable quantity, the absorptance as
and its average can be calculated using the transfer matrix method 72. For normal incidence, this quantity is independent of polarization.
@@
The size-dependence of for normal incidence at nm and nm are shown in Fig. 6(a). decreases with increasing film thickness for both wavelengths, with this decay occurring more rapidly at the longer wavelength. This trend can be explained by the distribution of the electric field within the metal film. Typically, the electric field is strongest near the surface of the metal film and exhibits an exponential decay as it penetrates deeper into the material. Thus, as the film thickness increases, the average field, hence, absorptance density decreases. This decay is more rapid for nm due to its shorter penetration depth ( nm) compared to the penetration depth of nm at nm as shown in Fig. 6 and Fig. 10(b).
Unlike the case of spheres, the size-dependence of the film temperature differs from . The temperature at the center of the illumination spot () as a function of film thickness is depicted in Fig. 6(b). Maximal temperature is reached at nm and nm, corresponding to the short and long wavelengths, respectively. This peak temperature can be explained from Fourier’s law of heat transfer (solution of the divergence in Eq. (10)), according to which, 222Solving the divergence in Eq. (10) simply along z-axis, we obtain . Since the temperature inside the film remains nearly constant along z-axis, and decreases linearly to at a distance outside the film, .. Thus, with increasing thickness, the temperature initially increases. However, due to the sharp decline of , with thickness, the temperature eventually decreases, resulting in a peak.
@@ 3.2.2 Determination of the PL
Fig. 7 presents the predicted PL obtained using Eq. (9) for Ag films of thickness ranging from nm to nm. For this calculation, we consider normal illumination at a wavelength of nm, with the emitted PL collected within a solid angle positioned directly above the illumination point. Calculations are performed at two incident intensities: one for which the films remain roughly at room temperature ( kW/cm2) and the other to elevate their temperature to K ( MW/cm2) as shown in Fig. 6(b). At room temperature, the PL at the three emission wavelengths ( nm, nm and nm) shows peaks at film thicknesses of nm, nm and nm, respectively. Similar to the case of the spheres, when the peak illumination intensity is increased to MW/cm2, the trends of SE remain unchanged so that Figs. 7(e)-(f) are hardly distinguishable from Figs. 7(b)-(c). However, the peak corresponding to aSE becomes narrower.
@@
Again, the calculations (14) closely match the results with the exact calculations (Eq. (9)), with the SE (Fig. 7(b)-(c),(e)-(f)) and low temperature aSE (Fig. 7(a)) being well approximated by for films up to nm, after which the approximation falls bellow the exact results 333For films thicker than nm, the PL calculated with Eq. (14) is less than that calculated with Eq. (9) as for .. Due to the multiplication by the increasing and the sharply decreasing (Fig. 6(a)), a peak appears in the PL. The stronger decay of at longer wavelengths causes these peaks to shift slightly towards smaller values of . However, at the higher temperatures, the aSE undergoes exponential amplification via the term, resulting in the sharper peak compared to that at the room temperature as illustrated in Fig. 7(d).
The PL calculated at the longer excitation wavelength of nm allows us to explore the PL behaviour deeper into the aSE regime, where the 2PA contribution takes over the 1PA contribution, see Figs. 16(a) and (d). For this excitation, while the PL in the near aSE region, (Fig. 16(b) and (e)) can be explained using the same reasoning applied to nm illustrated in Fig. 7(a) and (d), in the deeper aSE regime, the PL varies as (rather than ) at room temperature and as (rather than ) at the high temperature excitation.
Note that unlike the case of spheres, the PL from the layer decreases monotonically beyond the increase for small sizes. The reason for that is the absence of the (high-order) resonances that occur for the spheres, into which the configuration is tuning in and out.

### v2

@@ 3.1.2 Determination of the PL
As already noted above, the expression for PL (9) can be further simplified for a sphere. As discussed in Section 3.1.1, the temperature is fairly uniform in the sphere and the space-dependence of in Eq. (9) originates only from that of and , as indicated by Eq. (2.1). Further, we rely on the calculations of the mean free path of the non-thermal electrons 31 to justify the neglect of spatial broadening of the non-thermal electron distribution. These results motivate the use of the analytic result (2.1) also for non-uniform fields, i.e., to approximate the level of non-thermal electron population by simply summing over its local value across the particle volume. Accordingly, (5) can be expressed in terms of the spatially averaged, . With this consideration, (2.1) becomes independent of position. The volume integration over is now replaceable by the absorption cross-section of the sphere, which in turn, can be expressed as . Thus, Eq. (9) simplifies to
@@
Using this expression, we can determine the total PL solely via the absorption cross-section density evaluated at the pump frequency (, Fig. 2(b)) and emission frequency (, Fig. LABEL:Fig.alpha_vs_WL(a)), as well as the electron temperature (Fig. 2(c)). As seen in Fig. 3(a)-(c), the qualitative behaviour of the PL (Eq. (9)) and its approximation (Eq. (12)) is nearly identical except for the former being slightly higher for nm 222This can be shown analytically via the Mie solution; the proof of this behaviour for the case of layers is shown in footnote 4. .
@@
While the computational simplification associated with Eq. (12) is minor, its more significant contribution is to allow us to pinpoint the origins of the behaviour observed in Fig. 3. Specifically, the PL in Fig. 3(a)-(c) pertains roughly to room temperature for all sphere sizes, making the Bose function () essentially size-independent. In this scenario, the PL varies with sphere size as the product of (at the absorption wavelength), (at the emission wavelength) and the volume, . For spheres of radius nm (i.e., safely within the quasi-static regime), has a very weak dependence on the volume (see Fig. LABEL:Fig.ACS_vs_radius) so that the PL scales as . For spheres with sufficiently large radii with respect to the skin depth, the absorption is dominated by the surface layer of the particle, so that . Consequently, the PL scales linearly with (). For sphere sizes between these two extremes, the sub-volume scaling of the absorption and temperature, along with variations in with particle size owing to the sphere’s modal response (see Ref. [ 36] and Fig. 2(b)) result in the observed deviation from the volume scaling, namely, the peaks and troughs in the PL shown in all subplots of Fig. 3. Specifically, the non-monotonic behaviour in the PL observed for the larger sizes are due to the combined effect of the system tuning into and then out of the dipole resonance, and then tuning into the quadrupole mode. In that respect, considering the PL as being a volume effect for small spheres, and as a surface effect for larger spheres, is a crude description, which misses the dominant effect of the modal structure but captures the baseline.
@@
Fig. 3(d)-(f) shows the PL for a higher intensity which results in a size-dependent temperature, varying from room temperature to (see Fig. 2(c)). The higher varying temperature affects differently the different parts of the PL spectrum. First, it hardly affects the SE (Fig. 3(e)-(f)) which thus follows the same trend observed for the low-intensities, at least up to temperatures for which thermo-optic effects are still negligible 39, 67 (see Section 3.1.3 below). Indeed, for SE, the non-thermal contribution from the PA term dominates (see Eq. (2.1), Fig. 1 and Fig. LABEL:Fig.Ie_vs_radius_sphere), which is nearly temperature-independent since it originates from the non-thermal electron shoulders given in Eq. (5) (see SI Section S3 of Ref. [ 12]). In contrast, the aSE is exponentially sensitive to the temperature via the Bose function (see Fig. 1). As a result, the size-dependence of the PL is much stronger (see Fig. 3(d)), exhibiting superlinear scaling with the volume.
@@
Deeper into the aSE regime, the PA term becomes stronger than the 1PA term, see Fig. 1. To demonstrate this, we calculate the PL for a longer (off-resonant) excitation wavelength ( nm, i.e., smaller ), see Fig. LABEL:Fig.PL_900nm_sphere. In this case, while the near aSE (Fig. LABEL:Fig.PL_900nm_sphere(b) and (f)) grows as , for frequencies sufficiently deep into the aSE, the PL scales as at low intensity excitation (Fig. LABEL:Fig.PL_900nm_sphere(a)) and as at high intensity illumination (Fig. LABEL:Fig.PL_900nm_sphere(d)).
@@
Finally, in order to show the generality of our approach, and to demonstrate that the behaviour we identify is not qualitatively affected by the choice of the exciting laser frequency (e.g., its position with respect to resonance), we replicate the results of Fig. 3 (for which the illumination is on the red-side of the plasmon resonance) in Fig. LABEL:Fig.PL_488nm_n-1.8 and Fig. LABEL:Fig.PL_488nm_n-2.2 for backgrounds with higher refractive indices. This results in a red-shift of the resonances and introduces the higher-order modal response to the PL emission, particularly, for the larger spheres. However, the qualitative behaviour is similar to that with the lower background refractive index.
@@ 3.1.3 The thermo-optic effect
Illumination at high intensity can induce significant changes in the temperature of the spheres, consequently altering their permittivity via the so-called thermo-optic effect 68, 38, 69, and as a result, modify the absorptivity (hence, the emissivity) and finally the PL. To analyze this effect, we assume that the permittivity of the metal and the thermal conductivity of the surrounding medium depend linearly on the change of temperature, compute the thermo-derivative of the metal permittivity from ellipsometry data 70 and set the thermo-derivative of the thermal conductivity of the surrounding medium to Wm-1K-2, as in Ref. [ 67]; the temperature of the nanoparticle is then calculated using the method described by Un et al. 67.
Overall, as the intensity (hence, temperature) increases, the real part of the permittivity becomes more negative, while the imaginary part increases to higher positive values at most frequencies. This results in a blue shift of the resonance peak and a broadening of its spectral width, leading to enhanced absorption at off-resonance frequencies, see Fig. LABEL:Fig.abs_spec_TOE and Ref. [ 71, 72, 70, 38, 69]. As a result, slightly decreases at nm and slightly increases at nm, as seen in Fig. 2(b). The decrease in the former case leads to a reduction in the temperature for the hottest spheres (around nm), and hence, to the reduction of the aSE observed in Fig. 3(d). However, the SE PL is hardly affected by the temperature. Thus, the increase in at longer wavelengths results in a slight increase of the PL, as observed in Fig. 3(e) and (f). For excitation at nm, the temperature of the spheres (especially for nm) increases slightly compared to the value calculated for the linear case, leading to a slight increase in the PL for aSE, as shown in Fig. LABEL:Fig.PL_900nm_sphere(d)-(e). The thermo-optic increase of and leads to a slight increase in SE, as observed in Fig. LABEL:Fig.PL_900nm_sphere(f).
@@ 3.1.4 Comparison to single particle experiments
Our theoretical predictions can be compared to various single particle PL measurements. First, Fig. 4(a) shows that our calculations align with the experimental observations of the normalized SE PL (spectrally integrated over the range of nm to nm) measured from single gold nanospheres of different sizes, as reported by Gaiduk et al. (see Fig. S5 of Ref. [ 25]). The experimental results are presented as they appear in the referenced paper (in arbitrary units), and the corresponding theoretical calculations are scaled by a constant normalization factor to enable quantitative comparison with the experimental data. In this experiment, the illumination intensity was adjusted to keep the temperature rise relatively low (approximately below K). The PL is calculated using Eq. (9) with the refractive index of gold obtained from Ref. [ 73], and those of glass and glycerol being and , respectively. The normalized PL roughly scales with the volume until nm, but seems to saturate upon further increase in size; assuming this is the onset of a decrease, this is in line with our prediction.
@@
Size-dependence of the PL from larger particles was studied by Hu et al. 26 who demonstrated that the spectral peak of the PL from gold nano-disks on a SiO2-coated absorptive (silicon) substrate initially increases, reaches a maximum, and then decreases significantly with further increases in size. The absorption cross-section densities and the of the disks are calculated using COMSOL, considering normal incidence and emission perpendicular to the substrate, with the permittivity data for gold and Si taken from Ref. [ 73] and Ref. [ 74], respectively. Calculations based on Eq. (9) exhibit a qualitatively similar trend (even if at a slight spectral shift).
@@
Lastly, our theoretical prediction of PL are compared with the PL measurements by Lin et al. 28. This work showed the spectrally-integrated PL from single nanorods of different lengths with constant diameter along their main axis. In contrast to the nanospheres (and in similarity to films, shown later), the PL monotonically decreases with nanorod length. These results are reproduced in Fig. 4(c) along with the theoretical prediction, i.e., the spectral integration of Eq. (9). The optical response is calculated in COMSOL, where the substrate refractive index is and the permittivity of the silver rods is taken from Ref. [ 73]. The temperature is estimated with an approximate expression 65 and remains close to the ambient temperature (0.1% variation) for the laser illumination intensity considered ( mW/cm2). Good qualitative agreement is obtained for nanorods longer than nm, however, our prediction underestimates the PL for shorter nanorods, thus providing only a qualitative match for this regime (specifically, in a 30 nm wide wavelength range close to the pump). A similar observation was recently reported above the pump wavelength for Au rods 37; in that case, the discrepancy between theory and measurements was ascribed to interband transitions.
@@
Remarkably, while our theory (9) includes only emission events occurring in the conduction band, it successfully predicts the dependence of the emission on the particle size even in spectral regimes where interband emission events should occur (specifically, for the spheres and nano-disks (Fig. 4(a)-(b), respectively). This implies that, in these cases, interband transitions modify the emission features in a modest quantitative manner, but not qualitatively. We expect that inclusion of such transitions may yield a better quantitative match to the experimental results , due to a better match of the permittivity and non-thermal electron distribution.
3.2 Films under continuous illumination
@@
We now turn to study the PL from a thin film of a Drude metal using the approach used above. Specifically, we examine a thin silver film on a glass substrate illuminated at normal incidence (hence, no dependence on polarization , and , where is the component of the wavenumber k normal to the surface of the film) by a focused Gaussian CW beam with an intensity profile as depicted in Fig. 5(a).
@@ 3.2.1 Determination of the electron temperature
The absorbed power density and temperature distribution calculated by solving the heat equation (10) using the electromagnetic heat module of COMSOL Multiphysics are presented in Figs. 5(b)-(c), respectively. The absorbed power density decreases exponentially along the illumination direction (-axis) due to the short penetration depth (Fig. 5(d)). However, along the radius it exhibits nearly the same distribution as that of the incident intensity, as shown in Fig. 5(e).
@@
Generically, is uniform in the plane of the film (i.e., it is independent of and ) and varies only along the thickness of the film (i.e., it is only -dependent). The average value of along the thickness can be derived from the experimentally measurable quantity, the absorptance , as
and its average can also be calculated using the transfer matrix method 75. For normal incidence, this quantity is independent of polarization.
@@
The size-dependence of for normal incidence at nm and nm are shown in Fig. 6(a). decreases with increasing film thickness for both wavelengths, with this decay occurring more rapidly at the longer wavelength. This trend can be explained by the distribution of the electric field within the metal film. Typically, the electric field is strongest near the surface of the metal film and exhibits an exponential decay as it penetrates deeper into the material. Thus, as the film thickness increases, the average field, hence, absorptance density decreases. This decay is more rapid for nm due to its shorter penetration depth ( nm) compared to the penetration depth of nm at nm as shown in Fig. 6 and Fig. LABEL:Fig.alpha_vs_WL(b).
Unlike the case of spheres, the size-dependence of the film temperature differs from . The temperature at the center of the illumination spot () as a function of film thickness is depicted in Fig. 6(b). Maximal temperature is reached at nm and nm, corresponding to the short and long wavelengths, respectively. This peak temperature can be explained from Fourier’s law of heat transfer (solution of the divergence in Eq. (10)), according to which, 333Solving the divergence in Eq. (10) simply along z-axis, we obtain . Since the temperature inside the film remains nearly constant along z-axis, and decreases linearly to at a distance outside the film, .. Thus, with increasing thickness, the temperature initially increases. However, due to the sharp decline of , with thickness, the temperature eventually decreases, resulting in a peak.
@@ 3.2.2 Determination of the PL
Fig. 7 presents the predicted PL obtained using Eq. (9) for Ag films of thickness ranging from nm to nm. For this calculation, we consider normal illumination at a wavelength of nm, with the emitted PL collected within a solid angle positioned directly above the illumination point. As above, calculations are performed at two incident intensities: one for which the films remain roughly at room temperature () and the other to elevate their temperature to K ( MW/cm2) as shown in Fig. 6(b). At room temperature, the PL at the three emission wavelengths ( nm, nm and nm) shows peaks at film thicknesses of nm, nm and nm, respectively. Similar to the case of the spheres, when the peak illumination intensity is increased to MW/cm2, the trends of SE remain unchanged so that Figs. 7(e)-(f) are hardly distinguishable from Figs. 7(b)-(c). However, the peak corresponding to aSE becomes narrower.
@@
Again, the calculations (14) closely match the results with the exact calculations (Eq. (9)), with the SE (Fig. 7(b)-(c),(e)-(f)) and low temperature aSE (Fig. 7(a)) being well approximated by for films up to nm, after which the approximation falls below the exact results444For films thicker than nm, the PL calculated with Eq. (14) is less than that calculated with Eq. (9) because for ..
As for particles, we can now deduce the general scaling of the PL with the size from Eq. (14). Specifically, for very thin films ( nm), decays relatively slowly. As a result, due to the presence of the term in the PL expression in Eq. (14), the PL in this regime exhibits sublinear growth. For thicker films ( nm), due to the fact that total absorption becomes almost independent of film thickness (). Consequently, the PL scales as , as demonstrated in Fig. 7(c) and (f) and. In the intermediate region, a peak appears. The stronger decay of at longer wavelengths causes the PL peak to shift slightly towards smaller values of (see Fig. 7(d), (f) and Fig. LABEL:Fig_Film_Ag_900nm). As seen in Fig. 7(d), for higher temperatures, the aSE undergoes exponential amplification via the term, resulting in sharper peaks compared to those at the room temperature.
The PL calculated at the longer excitation wavelength of nm allows us to explore the PL behaviour deeper into the aSE regime, where the 2PA contribution takes over the 1PA contribution, see Figs. LABEL:Fig_Film_Ag_900nm(a) and (d). For this excitation wavelength, while the PL in the near aSE region (Fig. LABEL:Fig_Film_Ag_900nm(b) and (e)) can be explained using the same reasoning applied to nm illustrated in Fig. 7(a) and (d), in the deeper aSE regime, the PL varies as (rather than ) at room temperature and as (rather than ) at the high temperature excitation. It should be noted that, in the high-temperature 2PA aSE case shown in Fig. LABEL:Fig_Film_Ag_900nm(d), overestimates the exact calculation (9) due to the additional dependence of the exact PL on . As the temperature increases further, provides a closer estimate of the exact PL.
Note that unlike the case of particles, the PL from the film decreases monotonically beyond the increase for small sizes. The reason for that is the absence of the resonances that occur for particles, into which the system is tuning in and out with the varying size.